SELF- AND AIR-BROADENING AND SHIFT COEFFICIENTS OF CH_{4} LINES IN THE 3 μm REGION

V. MALATHY DEVI, D. CHRIS BENNER, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795; \underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences Division, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; G. GUELACHVILI, Laboratoire de Physique Moleculaire et Applications (LPMA), CNRS, Universite Paris-Sud, Batiment 350, 91405 Orsay-Cedex, France; and L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

Self- and air-broadening and shift coefficients for numerous CH_{4} transitions in the 3 μm region have been determined through analysis of high-resolution absorption spectra recorded by two different Fourier transform spectrometer (FTS) systems. The first set of spectra was recorded using the 1-m FTS at the McMath-Pierce facility of the National Solar Observatory at Kitt Peak; resolution was 0.01 cm^{−1}. This data set included both self-broadened CH_{4} spectra with cell path lengths of 5, 10, 25, 150, and 2500 cm and pressures ranging from 0.4 Torr to 456 Torr, and air-broadened spectra of ^{12}CH_{4} and ^{13}CH_{4} with cell path lengths of 5, 25, 50, and 150 cm and total pressures from about 100 Torr to 400 Torr. A second set of self-broadened CH_{4} spectra was recorded at higher resolution (0.0027 cm^{−1}) using the FTS of the Laboratoire de Physique Moleculaire et Applications in Orsay. Path lengths for these spectra varied from 1.7 cm to 4 m, and pressures ranged from less than 1 Torr to over 500 Torr. A multispectrum nonlinear least-squares procedure^{a} was used to fit over 40 spectra simultaneously to determine self- and air-broadening and shift coefficients for more than 500 lines; most of these are in the ν_{3} and ν_{2} + ν_{4} bands. Differences between air- and self-broadening and shifts of ^{12}CH_{4} lines will be discussed, and air-broadening and shift coefficients for ^{12}CH_{4} and ^{13}CH_{4} will be compared. We will also compare our measurements with the values on the current HITRAN compilation^{b} and with other available measurements.

aD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith, and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

bL. S. Rothman et al., JQSRT \underline{48}, 469-507 (1992); L. S. Rothman et al., JQSRT, in press (1998).